OpenStudy (karatechopper):

Can someone please check a trig function worksheet for me? I'm practicing! :)

OpenStudy (lfreeman1285):

I might be able to help!

OpenStudy (karatechopper):

Thanks! @IFreeman1285

OpenStudy (loser66):

WWWWWWWWWWWWWooooooooooooooooah!! you memorize all of them??? A/MA/Zing

OpenStudy (karatechopper):

Didn't memorize! I use the hand method. :)

OpenStudy (loser66):

teach me, please

OpenStudy (lfreeman1285):

memorizing them isn't incredibly difficult, I had to for precalculus. These look correct! I didn't find any mistakes in my quick glance. I'll continue looking though.

OpenStudy (karatechopper):

So basically each finger on the hand has a different angle. Thumb-90 Index-60 Middle-45 Ring-30 Thumb-0 The palm is always 2. On top= cosine (side where the thumb points up) On bottom=sine (side where the pinky finger is) Say for example we are given the problem find Cos(pi/6) We know that pi/6 is a 30 degree angle, thus we put our ring finger down. We also can tell from the unit circle that when cos is pi/6, cos is positive. We put our ring finger down and then see that we have up on the cosine side above the folded finger. We have 3 fingers (ON TOP-COSINE) Now, anytime we look at the fingers, we need to square root them. So we say sqrt of 3, and then we put the value above the 2 in the palm bc this isn't Tangent or Cotangent where we have to look at both Sine and Cosine. Therefore, the answer for cos(pi/6) is sqrt3/2. Make sense? :)

OpenStudy (karatechopper):

Thank you free man! Yeah this is a precalc review for a calc quiz I'm taking tomorrow haha.

OpenStudy (karatechopper):

@Loser66 there's an explanation example for the hand method I use.

OpenStudy (loser66):

Thank you.

OpenStudy (lfreeman1285):

it looks like you're going to do well on that quiz then haha

OpenStudy (karatechopper):

I would hope so! Are they all correct? The quiz tomorrow is timed and I have a lot of other tests/discussions, and work in school tomorrow. I hope I'm not worn out by the time I get to this quiz! :) .

Nnesha (nnesha):

tan 120 is it -sqrt{3} or positive ?

OpenStudy (lfreeman1285):

It's negative, because its in the second quadrant.

Nnesha (nnesha):

http://prntscr.com/8at9tn looks like square root tan isn't positive in 2nd q

Nnesha (nnesha):

no i mean on the worksheet :) there is just sqrt{3}

OpenStudy (lfreeman1285):

Tangents are negative in the second and fourth. Cosines are negative in the second and third. Sines are negative in the third and fourth.

OpenStudy (lfreeman1285):

So just being careful about which quadrant everything is in should help. Drawing the circle at the top of the test is what helped me.

Nnesha (nnesha):

yes i know.

Nnesha (nnesha):

that question was for kara :D

OpenStudy (karatechopper):

Thanks!!!

OpenStudy (karatechopper):

Help a lot guys!

OpenStudy (karatechopper):

Have a good day/night everyone! :) x

Nnesha (nnesha):

csc =1/sin so\[\huge\rm csc(\frac{ 7\pi }{ 6 })=\frac{ 1 }{ \frac{ -1 }{ 2 } }\]

OpenStudy (karatechopper):

-2!

Nnesha (nnesha):

yes there is -1/2 on the wrkshet

OpenStudy (karatechopper):

Lol yeah I just saw that haha. Fixed, thank you! :) x

Nnesha (nnesha):

11) sec 180 is wrong there is no way you can get 1 for sec tan can be one

Nnesha (nnesha):

good job :) 3 mistakes i guess typoz :D tan 120 sec 180 csc (7pi/6) :=)